# Proton Dynamics Scenarios in the Integrable Optics Test Accelerator (IOTA) at Fermilab

### v1

@@ Abstract
The Integrable Optics Test Accelerator (IOTA) at Fermilab provides a versatile platform for studying the interplay of space-charge, impedance, and non-linear optics in high-intensity hadron beams within synchrotrons and storage rings. This report examines the parameters and dynamics of 2.5 MeV proton beam operations in two configurations of the bare IOTA lattice: one for demonstrating Non-linear Integrable Optics with the Danilov-Nagaitsev magnet, and the other for use with electron cooling. We offer order-of-magnitude estimates of the transverse emittance growth rate as a function of beam intensity, highlighting contributions from residual gas scattering, intra-beam scattering, and space-charge effects. Under nominal conditions, the beam lifetime is projected to be less than 7 minutes at low intensity with the current vacuum quality, and fewer than 100,000 turns at high intensity due to strong space-charge effects. The calculations presented here will guide strategies to mitigate emittance growth and inform future IOTA experiments.
@@ 1 Introduction
Future proton drivers for neutron[1], neutrino[2], and muon generation[3] aim to deliver beams with intensity and phase-space density an order of magnitude greater than current achievements. At Fermilab, we have proposed the Accelerator Complex Evolution plan[4] to significantly increase the number of protons delivered to the Long-Baseline Neutrino Facility (LBNF) target compared to the PIP-II era[5]. A key component of this plan is replacing the Fermilab Booster[6], possibly with a new Rapid Cycling Synchrotron (RCS) capable of operating at high intensity and managing high space-charge tune shifts. To tackle emittance growth, beam loss, and instabilities in this high-intensity regime, it is crucial to advance our understanding of space-charge compensation, the interaction between coherent instabilities and space-charge[7, 8, 9], and the best practices for applying Landau damping while maintaining dynamic aperture. The proton program at the Integrable Optics Test Accelerator (IOTA)[10] employs 2.5 MeV (with $pc\approx 68.5$ MeV/c) protons to mimic beams with incoherent betatron tune-shifts approaching 0.5 in a controlled impedance environment, utilizing a digital wake-building feedback system known as the waker[11] (see Fig. 1 for the layout of the bare storage ring). IOTA enables a wide range of experiments on Non-linear Integrable Optics[12, 13], space-charge compensation[14], electron lens[15], instabilities, and Landau damping[16]. The knowledge gained will contribute to the design of future hadron synchrotrons and storage rings.
@@
While the low-energy proton beam provides a valuable platform for isolating the dynamics of intense space-charge from coherent instabilities, it is vulnerable to significant emittance growth and loss mechanisms. Single Coulomb scattering of protons by residual gas nuclei limits the beam lifetime through large-angle elastic scattering events, whereas multiple inelastic scattering with atomic electrons causes emittance growth. A 2.5 MeV proton beam with an rms emittance of 4 µm, circulating in IOTA with a nominal beta function of 4 m (maximum 10 m) and a chamber diameter of 50 mm, experiences emittance growth and beam loss on time scales of 70 seconds and 16 minutes, respectively, due to residual gas scattering at an effective residual gas density of $7.0\times 10^{8}$ cm-3, as measured in atomic hydrogen equivalent using the dynamics of single electrons[17] in IOTA. Additionally, intra-beam scattering (IBS) results in faster emittance growth, with a time scale of a few seconds at the phase-space densities required for studies involving intense space-charge. Therefore, sustaining the beam in IOTA will require compensating for heating mechanisms through cooling. In this report, we explore the expected range of emittance growth rates of the proton beam through various mechanisms, which will guide the commissioning of protons in the IOTA ring, procedures for planned experiments, and the design of the electron cooler[18] as part of the broader electron-lens research program.
@@ 2 Beam from the IOTA Proton Injector
Emittances ($\epsilon_{x},\epsilon_{y}$) 4.3, 3 3.09.03.0-9.0 µm
@@
Figure 2: Normalized rms emittance of protons as a function of beam current as measured in a previous test of the duoplasmatron source and the Low Energy Beam Transport section. Figure reproduced from Wai-Ming Tam’s PhD thesis[23].
@@ 3 Linear Optics in IOTA
A wide range of linear optics configurations can be achieved in IOTA by adjusting the field strength of individual quadrupoles. A total of 79 free parameters allow us to control dispersion, one-turn phase advance, momentum compaction, and transverse coupling. Figure 3 shows examples of betatron amplitudes ($\beta_{x,y}$) and dispersion ($D_{x,y}$) as functions of position ($s$) around the ring, computed using the MAD-X[24] code, for two lattice configurations. The black curves represent the optics functions used for experiments with the Danilov-Nagaitsev (DN) non-linear magnet[25], the blue curves depict an ensemble of possible configurations for electron cooling (ECOOL), and the orange curves indicate a reference lattice from this ensemble. A separate document[26] details the design of the electron cooling lattices. The transverse acceptances ($\epsilon_{mx,y}\equiv\max{r_{x,y}^{2}/\beta_{x,y}}$) of the lattice quantify the maximum phase-space area the beam can occupy and are dependent on the aperture radii ($r_{x,y}$). The non-linear magnets in sections BR and BL, as well as the planned Ionization Profile Monitor[27] in section E, impose strict aperture limitations on the beam, significantly reducing the acceptance of the lattice. Table 2 highlights this effect by comparing the acceptances with aperture restrictions to a configuration where the aperture is the nominal beam pipe radius of 25 mm throughout the ring. The transverse optics also influence the longitudinal dynamics of the ring.
@@
where $\eta$ is the phase slip factor, and $n$ is the number of standard deviations of the injected momentum spread to be captured. $\beta$ and $\gamma$ are the relativistic velocity and energy parameters, respectively. The maximum value of rf voltage is constrained by the available rf power driving the cavity, which is estimated to be 1 kV. In the bunched beam configuration, the steady-state rf voltage $V$ and the phase slip factor determine the synchrotron tune $Q_{s}$, equilibrium momentum spread $\sigma_{\delta,\mathrm{eq}}$, and bunch length $\sigma_{s}$ as follows,[30]
@@
where $V_{1}$ is the final rf voltage in equilibrium, and $t_{1}$ is the number of turns used to ramp the rf voltage. The starting value of voltage $V_{0}$, seen by the beam right after injection, must be non-zero and is given by,
@@
where $Q_{s}$ is the equilibrium synchrotron tune corresponding to $V_{1}$, evaluated using Eq. (2a), and $n_{\mathrm{ad}}$ is the adiabaticity factor. The results of numerical simulations conducted using the PyORBIT code[32, 33] are shown in Fig. 4. In these simulations, we initialize a coasting beam modeled using 105 macro-particles, with the nominal energy and transverse emittances specified in Table 1, and track the beam for 2000 turns in the DN lattice. The adiabatic capture parameters are set to $t_{1}=1000$ and $n_{\mathrm{ad}}=10$. We repeat these simulations for multiple values of injected momentum spread, beam energy, and final voltage, within the ranges given in Table 1. Each gray band in Fig. 4 depicts $\sigma_{\delta,\mathrm{eq}}$ as a function of $V_{1}$ starting from a fixed value of $\sigma_{\delta,\mathrm{inj}}$. The lower and upper limits of the band corresponds to beam energies 2.5 MeV and 2.3 MeV respectively. The blue shaded region includes all configurations where $V_{1}\geq V_{\mathrm{min}}(\sigma_{\delta,\mathrm{inj}})$, where we can calculate the minimum capture voltage using Eq. (1). Both the DN lattice and the ensemble of ECOOL lattices exhibit similar values of final momentum spread since the phase-slip factor for all lattices is $|\eta|\sim 1$. Consequently, the range of equilibrium momentum spread shown in Table 2 is similar for both the DN and the ECOOL lattices.
@@
where $\sigma_{x,y}(s)$, $\mathrm{d}N/\mathrm{d}s$ are the transverse beam sizes, peak line density of the beam, and $r_{p}$ is the classical radius of the proton. The incoherent tune shift evolves with time as the phase space of the beam changes due to various heating mechanisms. Hence, we estimate the tune shifts only at the initial conditions of the beam. In the case of coasting beams, we use the values of injected transverse emittances and momentum spread listed in Table 1, while in the case of bunched beams, we keep the same value of transverse emittance but use the values of relative momentum spread and bunch length after adiabatic capture. Table 2 shows the range of beam intensity required to reach the intense space-charge regime for two lattices, taking into account the full range of energy, emittance, and momentum spread. The primary challenge to actually realizing such high tune shifts in the low-energy regime is the presence of various mechanisms that inflate the phase space and limit the storage time.
@@ 4 Emittance growth and beam lifetime due to residual gas scattering
Multiple scattering events increase the angular distribution of the beam, leading to steady growth of transverse emittance ($\mathrm{d}\epsilon_{x,y}/\mathrm{d}t$), given by (Combining Eq. (10) and (11) from Sec. 3.3.2 in [35]),
@@ 5 Emittance growth due to intra-beam scattering
Elastic scattering between protons inside the beam leads to momentum transfer between different degrees of freedom and an overall decrease in the phase-space density. The growth rates for transverse emittance and relative momentum spread due to intra-beam scattering (IBS) in a smooth round beam stored in a ring with zero dispersion can be approximated as,[37]
@@ 6 Space-Charge Dynamics in bare IOTA
Apart from short-range interactions due to elastic scattering, collective effects driven by space-charge and wakefields contribute to beam dynamics at high intensities. In the case of IOTA proton operations, wakefields generated by the interaction of the beam with the vacuum chamber are negligible, while space-charge effects dominate. The transverse rms beam size ($\sigma_{x,y}$) relative to the Debye length ($\lambda_{D,p}$) of the beam core dictates whether coherent effects or single-particle dynamics govern the evolution of phase-space. We define the Debye length in the center-of-mass frame, at the core of a Gaussian beam as follows,
@@
where $n_{p}$, $v_{\mathrm{th}}$ and $\omega_{p}=\sqrt{n_{p}e^{2}/(m_{p}\epsilon_{0})}$ are the number density, rms thermal velocity, and plasma frequency of protons at the beam core, respectively. In addition, we can define a characteristic length scale of separation between protons in the beam frame as $d_{p}\equiv(4\pi n_{p}/3)^{-1/3}$ and a length scale associated with the period of plasma oscillation $l_{p}\equiv 2\pi c\beta\gamma/\omega_{p}$ in the lab frame. The range of the characteristic length scales, for both coasting and bunched beams at the incoherent tune shift $|\delta\nu_{y,\mathrm{SC}}|=0.5$, in the DN and ECOOL lattices over the full range of beam parameters, including energy, emittances, momentum spread and bunch length are shown in Table 4. Based on these ranges, we can conclude that the Debye length is comparable or larger than the transverse size of the bunch. Thus, single-particle resonances such as those from magnetic field multipoles or from space-charge modulations (structure resonances) drive emittance growth and beam loss.[41] We use the Particle-in-Cell (PIC) simulation code PyORBIT to model the evolution of phase-space in the IOTA lattice.
@@
Transverse grid size ($\Delta x_{\mathrm{PIC}}$) $50/128\approx 0.39$ $m$
@@
We choose the 2.5 D PIC space charge model in PyORBIT, which applies transverse kicks to all macro-particles at various positions in the lattice, assuming that the number density profile of individual transverse slices of the bunch remain constant over its longitudinal extent. As listed in Table 4, the transverse grid size ($\Delta x_{\mathrm{PIC}}$) is fixed, and is consistent with the requirement that it must be larger than the characteristic distance between individual protons, while being smaller than the Debye length and the beam size. Conversely, the longitudinal grid has a fixed resolution while the extent is dynamically adapted to the positions of the macro-particles. The maximum distance between space-charge kicks ($\Delta s_{\mathrm{PIC}}$) is much smaller than the characteristic length of a plasma oscillation. Further, we choose the number of macro-particles to exceed the total number of cells in the 2 D grid, and thus obtain rms values of phase-space variables with relative uncertainties at the few-percent level. To be consistent with the calculations in the previous sections, we use elements of bare IOTA (only dipoles and quadrupoles) with no errors, no multipole fields and use a uniform aperture diameter of 50 mm everywhere. We initialize the beam with a Gaussian distribution in transverse phase-space and either a uniform or Gaussian distribution in longitudinal space depending on whether the beam is coasting or bunched, respectively.
@@
Figure 9: Emittance growth of coasting beam driven by betatron mismatch due to space-charge in the DN lattice. Panels (a) and (b) show the horizontal and vertical emittance growth, respectively, in IOTA over 20 turns after injection. The solid lines represent emittance growth when starting with Twiss parameters of the bare lattice, while the dashed line shows the same when starting with optimized Twiss parameters. Panel (c) shows the difference of optimum Twiss parameters from the bare values as a function of beam current.
@@
As the injected beam relaxes into an equilibrium distribution over the first few turns, the mechanism of initial emittance growth is the mismatch between the initial beam distribution and the effective focusing created by the lattice and the beam’s self-field. To minimize this mismatch-driven emittance growth, we optimized the Twiss parameters of the injected distribution. Specifically, we defined a loss function based on the transverse emittance growth after 20 simulated turns and used the Broyden–Fletcher–Goldfarb–Shanno algorithm to search over $\beta_{x}$, $\beta_{y}$, and $\eta_{x}$ (with all other Twiss functions set to zero since the lattice is mirro-symmetric about the center). Figure 9 shows the results of this optimization for coasting-beam injection into the DN lattice as a function of beam current. As seen in panels (a) and (b), the optimization yields up to a 10 % reduction in final transverse emittance at 10 mA, with smaller gains at lower currents. The optimal Twiss values in panel (c) may serve as a guide when commissioning IOTA.
@@
We illustrate the long-term evolution of bunched beams at four intensities[47] in the DN lattice, using four simulation models: a 2.5 D transverse PIC model in PyORBIT, a similar 2.5 D PIC model in XSuite[48], a frozen model in XSuite, and an adaptive-frozen model[49] in MAD-X. [^3] Both frozen models are implementations of the Bassetti-Erskine formula[46] which calculates transverse fields for a charge distribution which is Gaussian in transverse and uniform in longitudinal. Figure 13 compares rms emittance and beam loss, while Fig. 14 compares the transverse profiles after $10^{4}$ turns. At a current of 1 µA, the transverse incoherent tune shift due to space charge is $\sim 2\times 10^{-4}$, leading to linear dynamics where the phase-space distribution remains Gaussian and emittance growth is minimal, consistent across all three codes. At moderate beam currents corresponding to an incoherent tune shift of less than 0.3, significant increases in transverse rms emittance are observed in all models, albeit with some quantitative differences. While the emittance evolution in the self-consistent PIC models in XSuite and PyORBIT are similar, the transverse beam distributions and the beam loss are different. Specifically, the results from the XSuite PIC model indicates the formation of more intense halo as seen in Fig. 14 compared to the PyORBIT model, and hence more beam loss. Once the incoherent tune shift exceeds 0.3, particles at the center of the beam cross the integer resonance, resulting in rapid beam blow-up, with all four simulations diverging in emittance growth and beam loss. These simulations neglect realistic magnetic field errors and multipole fields which can excite further resonances and lead to faster emittance growth and beam loss.

### v2

@@ Abstract
The Integrable Optics Test Accelerator (IOTA) at Fermilab provides a versatile platform for studying the interplay of space-charge, impedance, and non-linear optics in high-intensity hadron beams within synchrotrons and storage rings. This report examines the parameters and dynamics of 2.5 MeV proton beam operations in two configurations of the bare IOTA lattice (dipoles, quadrupoles, sextupoles, and rf cavity only): one for demonstrating Non-linear Integrable Optics with the Danilov-Nagaitsev magnet, and the other for use with electron cooling. We offer order-of-magnitude estimates of the transverse emittance growth rate as a function of beam intensity, highlighting contributions from residual gas scattering, intra-beam scattering, and space-charge effects. Under nominal conditions, the beam lifetime is projected to be less than 7 minutes at low intensity with the current vacuum quality, and fewer than 100,000 turns at high intensity due to strong space-charge effects. The calculations presented here will guide strategies to mitigate emittance growth and inform future IOTA experiments.
@@ 1 Introduction
Future proton drivers for neutron [1], neutrino [2], and muon generation [3] aim to deliver beams with intensity and phase-space density an order of magnitude greater than current achievements. At Fermilab, we have proposed the Accelerator Complex Evolution plan [4] to significantly increase the number of protons delivered to the Long-Baseline Neutrino Facility (LBNF) target compared to what is expected from the PIP-II upgrade [5]. A key component of this plan is replacing the Fermilab Booster [6], possibly with a new Rapid Cycling Synchrotron (RCS) capable of operating at high intensity and managing high space-charge tune shifts. To tackle emittance growth, beam loss, and instabilities in this high-intensity regime, it is crucial to advance our understanding of space-charge compensation, the interaction between coherent instabilities and space-charge [7, 8, 9], and the best practices for applying Landau damping while maintaining dynamic aperture. The proton program at the Integrable Optics Test Accelerator (IOTA) [10] employs 2.5 MeV (with $pc\approx 68.5$ MeV/c) protons to mimic beams with incoherent betatron tune-shifts approaching 0.5 in a controlled impedance environment, utilizing a digital wake-building feedback system known as the waker [11] (see Fig. 1 for the layout of the bare storage ring including dipoles, quadrupoles, sextupoles, and the rf cavity). IOTA enables a wide range of experiments on Non-linear Integrable Optics [12, 13], space-charge compensation [14], electron lens [15], instabilities, and Landau damping [16]. The knowledge gained will contribute to the design of future hadron synchrotrons and storage rings.
@@
While the low-energy proton beam will provide a valuable platform for isolating the dynamics of intense space-charge from coherent instabilities, it will be vulnerable to significant emittance growth and loss mechanisms. Estimates of single Coulomb scattering of protons by residual gas nuclei suggest that beam lifetime will be limited through large-angle scattering events, whereas multiple inelastic scattering with atomic electrons will cause emittance growth. A 2.5 MeV proton beam with an rms emittance of 4 µm, circulating in IOTA with a nominal beta function of 4 m (maximum 10 m) and a chamber diameter of 50 mm, will experience emittance growth and beam loss on time scales of 70 seconds and 16 minutes, respectively, due to residual gas scattering at an effective residual gas density of $7.0\times 10^{8}$ cm-3, as measured in atomic hydrogen equivalent using the dynamics of single electrons [17] in IOTA. Additionally, intra-beam scattering (IBS) will result in faster emittance growth, with a time scale of a few seconds at the phase-space densities required for studies involving intense space-charge. Therefore, sustaining the beam in IOTA will require compensating for heating mechanisms through cooling. In this report, we explore the expected range of emittance growth rates of the proton beam through various mechanisms, which will guide the commissioning of protons in the IOTA ring, procedures for planned experiments, and the design of the electron cooler [18] as part of the broader electron-lens research program.
@@ 2 Beam from the IOTA Proton Injector
Geometric emittances ($\epsilon_{x},\epsilon_{y}$) 4.3, 3 3.09.03.0-9.0 µm
@@
Figure 2: Normalized rms emittance of protons as a function of beam current as measured in a previous test of the duoplasmatron source and the Low Energy Beam Transport section Reproduced with permission from [23].
@@ 3 Linear Optics in IOTA
A wide range of linear optics configurations can be achieved in IOTA by adjusting the field strength of individual quadrupoles. A total of 79 free parameters allow us to control dispersion, one-turn phase advance, momentum compaction, and transverse coupling. Figure 3 shows examples of betatron amplitudes ($\beta_{x,y}$) and dispersion ($D_{x,y}$) as functions of position ($s$) around the ring, computed using the MAD-X [24] code, for two lattice configurations. The black curves represent the optics functions used for experiments with the Danilov-Nagaitsev (DN) non-linear magnet [25], the blue curves depict an ensemble of possible configurations for electron cooling (ECOOL), and the orange curves indicate a reference lattice from this ensemble. The DN lattice has already been implemented for 150 MeV electron operations in IOTA [26, 27] and benchmarked with linear optics calculations in MAD-X. A separate document [28] details the design of the electron cooling lattices for future use. The transverse acceptances ($\epsilon_{mx,y}\equiv\min{r_{x,y}^{2}/\beta_{x,y}}$) of the lattice quantify the maximum phase-space area the beam can occupy and are dependent on the aperture radii ($r_{x,y}$). The non-linear magnets in sections BR and BL, as well as the planned Ionization Profile Monitor [29] in section E, impose strict aperture limitations on the beam, significantly reducing the acceptance of the lattice. Table 2 highlights this effect by comparing the acceptances with aperture restrictions to a configuration where the aperture is the nominal beam pipe radius of 25 mm throughout the ring. The transverse optics also influence the longitudinal dynamics of the ring.
@@
where $\eta$ is the phase slip factor, and $n$ is the number of standard deviations of the injected momentum spread to be captured. $\beta$ and $\gamma$ are the relativistic velocity and energy parameters, respectively. The maximum value of rf voltage is constrained by the available rf power driving the cavity, which is estimated to be 1 kV. In the bunched beam configuration, the steady-state rf voltage $V$ and the phase slip factor determine the synchrotron tune $Q_{s}$, and the equilibrium momentum spread $\sigma_{\delta,\mathrm{eq}}$, which in turn determines the bunch length $\sigma_{s}$ as follows, [32]
@@
where $V_{1}$ is the final rf voltage in equilibrium, and $t_{1}$ is the number of turns used to ramp the rf voltage. The starting voltage $V_{0}$ must be non‑zero in order to keep the instantaneous synchrotron frequency finite and to satisfy the adiabaticity requirement (the relative change of the synchrotron‑bucket area is kept much smaller than the synchrotron frequency throughout the ramp). Its value is determined by the chosen adiabaticity factor $n_{\mathrm{ad}}$ and can be made arbitrarily small:
@@
where $Q_{s}$ is the equilibrium synchrotron tune corresponding to $V_{1}$, evaluated using Eq. (2a). The results of numerical simulations conducted using the PyORBIT code [34, 35] are shown in Fig. 4. In these simulations, we initialize a coasting beam modeled using 105 macro-particles, with the nominal energy and transverse emittances specified in Table 1, and track the beam for 2000 turns in the DN lattice, accounting for single-particle dynamics only. The adiabatic capture parameters are set to $t_{1}=1000$ and $n_{\mathrm{ad}}=10$. We repeat these simulations for multiple values of injected momentum spread, beam energy, and final voltage, within the ranges given in Table 1. Each gray band in Fig. 4 depicts $\sigma_{\delta,\mathrm{eq}}$ as a function of $V_{1}$ starting from a fixed value of $\sigma_{\delta,\mathrm{inj}}$. The lower and upper limits of the band corresponds to beam energies 2.5 MeV and 2.3 MeV respectively. The blue shaded region includes all configurations where $V_{1}\geq V_{\mathrm{min}}(2\sigma_{\delta,\mathrm{inj}})$, where the minimum voltage is calculated using Eq. (1) in order to capture twice the injected rms momentum spread. Since the initial momentum distribution is Gaussian, beam loss is less than 5 % inside the blue shaded region. Both the DN lattice and the ensemble of ECOOL lattices exhibit similar values of final momentum spread since the phase-slip factor for all lattices is $|\eta|\sim 1$. Consequently, the range of equilibrium momentum spread shown in Table 2 is similar for both the DN and the ECOOL lattices.
@@
where $\sigma_{x,y}(s)$, $\mathrm{d}N/\mathrm{d}s$ are the transverse beam sizes, peak line density of the beam, and $r_{p}$ is the classical radius of the proton. The incoherent tune shift generally reduces with time as the emittance of the beam grows due to various heating mechanisms. For instance, doubling the transverse emittance leads to the incoherent tune shift reducing by a factor of 2 as a consequence. Here, we estimate the tune shifts only at the initial conditions of the beam. In the case of coasting beams, we use the values of injected transverse emittances and momentum spread listed in Table 1, while in the case of bunched beams, we keep the same value of transverse emittance but use the values of relative momentum spread and bunch length after adiabatic capture. Table 2 shows the range of beam intensity required to reach the intense space-charge regime for two lattices, taking into account the full range of energy, emittance, and momentum spread. The primary challenge to actually realizing such high tune shifts in the low-energy regime is the presence of various mechanisms that inflate the phase space and limit the storage time.
@@ 4 Emittance growth and beam lifetime due to residual gas scattering
Multiple inelastic scattering events with atomic electrons increase the angular distribution of the beam, leading to steady growth of transverse emittance ($\mathrm{d}\epsilon_{x,y}/\mathrm{d}t$), given by (Combining Eq. (10) and (11) from Sec. 3.3.2 in [37]),
@@ 5 Emittance growth due to intra-beam scattering
Elastic scattering between protons inside the beam leads to momentum transfer between different degrees of freedom and an overall decrease in the phase-space density. To estimate the magnitude of this effect, we employ a simple model for the growth rates of the transverse emittance and the relative momentum spread caused by intra‑beam scattering (IBS). Within the smooth‑lattice approximation—assuming zero dispersion and a round beam—the growth rates can be expressed as, [39]
@@ 6 Space-Charge Dynamics in bare IOTA
Apart from short-range interactions due to elastic scattering, collective effects driven by space-charge and wakefields contribute to beam dynamics at high intensities. In the case of IOTA proton operations, wakefields generated by the interaction of the beam with the vacuum chamber are negligible, while space-charge effects dominate. The transverse rms beam size ($\sigma_{x,y}$) relative to the Debye length ($\lambda_{D,p}$) of the beam core dictates whether coherent effects or incoherent single-particle effects govern the evolution of phase-space. We define the Debye length in the center-of-mass frame, at the core of a Gaussian beam as follows,
@@
where $n_{p}$, $v_{\mathrm{th}}$ and $\omega_{p}=\sqrt{n_{p}e^{2}/(m_{p}\epsilon_{0})}$ are the number density, rms thermal velocity, and plasma frequency of protons at the beam core, respectively. In addition, we can define a characteristic length scale of separation between protons in the beam frame as $d_{p}\equiv(4\pi n_{p}/3)^{-1/3}$ and a length scale associated with the period of plasma oscillation $l_{p}\equiv 2\pi c\beta\gamma/\omega_{p}$ in the lab frame, both at the core of the beam. The range of the characteristic length scales, for both coasting and bunched beams at the incoherent tune shift $|\delta\nu_{y,\mathrm{SC}}|=0.5$, in the DN and ECOOL lattices over the full range of beam parameters, including energy, emittances, momentum spread and bunch length are shown in Table 4. Based on these ranges, we can conclude that the Debye length is comparable or larger than the transverse size of the bunch. Thus, single-particle resonances such as those from magnetic field multipoles or from space-charge modulations (structure resonances) drive emittance growth and beam loss. [43] We use the Particle-in-Cell (PIC) simulation code PyORBIT to model the evolution of phase-space in the IOTA lattice.
@@
Transverse grid size ($\Delta x_{\mathrm{PIC}}=\Delta y_{\mathrm{PIC}}$) $50/128\approx 0.39$ $m$
@@
We choose the 2.5 D PIC space charge model in PyORBIT, which applies transverse kicks to all macro-particles at various positions in the lattice, assuming that the number density profile of individual transverse slices of the bunch remain constant over its longitudinal extent. Table 4 shows that the transverse grid spacing is set to the pipe diameter divided by the chosen grid resolution. This choice satisfies two constraints: the cell size is larger than the typical inter‑proton spacing, yet smaller than both the Debye length and the overall beam size. With a total of $10^{5}$ macro-particles spread over a transverse grid of $128^{2}$, cells near the beam core have at least 6 macro-particles each. Conversely, the longitudinal grid has a fixed resolution while the extent is dynamically adapted to the positions of the macro-particles. The maximum distance between space-charge kicks ($\Delta s_{\mathrm{PIC}}$) is much smaller than the characteristic length of a plasma oscillation. Further, we choose the number of macro-particles to exceed the total number of cells in the 2 D grid, and thus obtain rms values of phase-space variables with relative uncertainties at the few-percent level. To be consistent with the calculations in the previous sections, we use linear elements of bare IOTA (only dipoles and quadrupoles) with no errors, no multipole fields, an uniform aperture diameter of 50 mm everywhere, and open boundary conditions for the space-charge field. We initialize the beam with a Gaussian distribution in transverse phase-space and either a uniform or Gaussian distribution in longitudinal space depending on whether the beam is coasting or bunched, respectively.
@@
Figure 9: Emittance growth of coasting beam driven by betatron mismatch due to space-charge in the DN lattice. Panels (a) and (b) show the horizontal and vertical emittance growth, respectively, in IOTA over 20 turns after injection. The solid lines represent emittance growth when starting with Twiss parameters of the bare lattice (including linear elements without space-charge de-focusing), while the dashed line shows the same when starting with optimized Twiss parameters. Panel (c) shows the difference of optimum Twiss parameters from the bare values as a function of beam current.
@@
As the injected beam relaxes into an equilibrium distribution over the first few turns, the mechanism of initial emittance growth is the mismatch between the initial beam distribution and the effective focusing created by the lattice and the beam’s self-field. To minimize this mismatch-driven emittance growth, we optimized the Twiss parameters of the injected distribution. Specifically, we defined a loss function based on the transverse emittance growth after 20 simulated turns and used the Broyden–Fletcher–Goldfarb–Shanno algorithm to search over $\beta_{x}$, $\beta_{y}$, and $\eta_{x}$ (with all other Twiss functions set to zero since the lattice is mirror-symmetric about the center). Figure 9 shows the results of this optimization for coasting-beam injection into the DN lattice as a function of beam current. As seen in panels (a) and (b), the optimization yields up to a 10 % reduction in final transverse emittance at 10 mA, with smaller gains at lower currents. The optimal Twiss values in panel (c) may serve as a guide when commissioning IOTA.
@@
We illustrate the long-term evolution of bunched beams at four intensities [49] in the DN lattice, using four simulation models: a 2.5 D transverse PIC model in PyORBIT, a similar 2.5 D PIC model in XSuite [50], a frozen model in XSuite, and an adaptive-frozen model [51] in MAD-X. [^3] Both frozen models represent the three‑dimensional charge distribution as a transverse 2‑D Gaussian multiplied by a longitudinal 1‑D Gaussian. The transverse electric field of the 2‑D Gaussian is obtained from the Bassetti‑Erskine formula [48] and then weighted by the local longitudinal charge density to calculate the transverse kicks. Both frozen models separate the 3D spatial distribution into a transverse 2D Gaussian weighted with a longitudinal 1D Gaussian. Figure 13 compares rms emittance and beam loss, while Fig. 14 compares the transverse profiles after $10^{4}$ turns. At a current of 1 µA, the transverse incoherent tune shift due to space charge is $\sim 2\times 10^{-4}$, leading to linear dynamics where the phase-space distribution remains Gaussian and emittance growth is minimal, consistent across all four models. At moderate beam currents corresponding to an incoherent tune shift of less than 0.3, significant increases in transverse rms emittance are observed in all models, albeit with some quantitative differences.
While the emittance evolution in the self-consistent PIC models in XSuite and PyORBIT are similar, the transverse beam distributions and the beam loss are different. Specifically, the results from the XSuite PIC model indicates the formation of more intense halo as seen in Fig. 14 compared to the PyORBIT model, and hence more beam loss. This discrepancy can be traced to two implementation differences: (i) PyORBIT is configured to use a fixed transverse grid, whereas the XSuite PIC model adapts the grid to the instantaneous rms beam size; and (ii) the number and longitudinal positions of the space‑charge kicks differ between the codes (both sets were individually converged, but the discretization is not identical). The frozen calculations also diverge. In the XSuite frozen model the transverse emittance is assumed constant, so the transverse beam size remains fixed with time even as the actual emittance grows; consequently the space‑charge force does not weaken as the beam blows up, leading to a faster emittance rise. The MAD‑X adaptive‑frozen model updates the beam size (and thus the space‑charge fields) turn‑by‑turn, which reduces the field strength as the beam expands and moderates the emittance growth. Moreover, MAD‑X applies both transverse and longitudinal apertures at the exit of every element, while XSuite is configured to impose a single transverse aperture at the ring entrance. Once the incoherent tune shift exceeds 0.3, particles at the center of the beam cross the integer resonance, resulting in rapid beam blow-up, with all four simulations diverging in emittance growth and beam loss.
The above discussion provides a starting point for understanding these differences, and a detailed benchmarking study of the various models using the IOTA 2.5 MeV proton‑beam dynamics scenario will be presented in future work (see Ref. [51] for previous benchmarking efforts involving MAD‑X and PyORBIT). These simulations also neglect realistic magnetic field errors and multipole fields which can excite further resonances and lead to faster emittance growth and beam loss.